OpenStudy (anonymous):

If the odometer on the car reads 25 km at the beginning of a trip and 65 km a half hour later, what is the average speed of the car?

OpenStudy (anonymous):

Add the speeds, then divide by 2.

OpenStudy (gebooors):

Distance driven during half an hour is 65 - 25 km. \[v = s/t = 40 km / 0,5 h\]

OpenStudy (anonymous):

I must be brain dead haha.

OpenStudy (gebooors):

Just look for units, no need any conversion in time and distance.

OpenStudy (anonymous):

80 km/hour